□(Round to four decimal places as needed) DRoundtofour decinal places as needed) O Type here to
Answer #1

M=26

S=0.06

A)X1=25.9

Z1=X1-M/S

=25.9-26/0.06

=-1.67

P(X1<25.9)

=P(Z1<-1.67)

=0.5-0.4525

=0.0475. Or 4.75%

B)X1=25.86. X2=26.14

Z1=x1-M/s

=25.86-26/0.06

=-2.33

Z2=x2-M/s

=26.14-26/0.06

=2.33

P(-2.33<z<2.33)

=0.4909+0.4909

=0.9818. Or 98.18%

C) probability rod discarded =0.0091 X 2 =0.0182

E(x)=np

=5000*0.0182

=91

d)probability to manufacture rod between 25.9 and 26.1

=0.4525x2

=0.905

E(x)=np

=1000x0.905

=9050